A pipeline contractor can purchase a needed truck for $42,000. Its estimated life is 6 years, and it has no salvage value. Maintenance is estimated to be $2,500 per year. Operating expense is $60 per day. The contractor can hire a similar unit for $170 per day. MARR is 7%.

How many days per year must the truck’s services be needed such that the two alternatives are equally costly?

Determine the dollar amount of annual savings generated by using the preferred alternative rather than the nonpreferred.
Answer #1

Let the no. of days of operation be n, then

Annual cost of purchase = 42000 * (A/P, 7%,6) + 2500 + 60 * n

Annual cost of hire = 170 * n

For breakeven

42000 * (A/P, 7%,6) + 2500 + 60 * n = 170 * n

42000 * 0.209796 + 2500 = 170 * n - 60 * n

11311.42 = 110 * n

n = 11311.42 / 110 = 102.83 ~ 103 days